OpenStudy (anonymous):

How is x squared + 3c - ( x squared - 3x)= 6x

OpenStudy (anonymous):

plus 3x**

OpenStudy (anonymous):

\(x^2 +3c-(x^2 +3x)=6x\) are you solving for x, and c is a constant?

OpenStudy (anonymous):

\[x^2 + 3x - ( x^2 -3x ) = 6x\] (opening bracket and multiplying with -ve) \[x^2 + 3x - x^2 + 3x =6x \] (x^2 - x^2 =0 ) \[6x = 6x\]

OpenStudy (anonymous):

ah. so c is x...............

OpenStudy (anonymous):

exactly;)

OpenStudy (anonymous):

wait. that means it's solve for c. lol